calculate the mass of hcl in grams when 250cm3 of water is added to 100cm3 of 0.5M hcl
1
Expert's answer
2020-05-01T14:23:14-0400

"M1\times V1=M2\times V2"

"0.5\times0.1=M2\times0.35"

"0.05=M2\times0.35"

M2=0.14

"n=M\times V=0.14\times0.35=0.05"


"m=n\times M=0.05\times36.4610=1.82" g